OpenStudy (anonymous):

factor the difference of the two squares x^8-y^10

OpenStudy (anonymous):

(x^4 + y^5)(x^4 - y^5) Note: y^10 is a perfect square becuase y^5 x y^5 = y^10 also x^8 is a perfect square becuase x^4 x x^4 = x^8